Find e At where 2 5 A - 2 -4
Answer #1

V (22) 5-1 -4-2 al det(A-1)=0 & 12-11-4-1) + 2/5 -8 + 4d-2y + y +19=9 2x + 1? +2=0 da lla -iti .2 Įda -2754212 1.e da -It is